A company plans to make 10 liters of a 17% acid solution by mixing a 20% acid solution and a 15% acid solution. How much of each solution should be used.
Answer by ewatrrr(24785)


Hi,
Let x represent the amount of the more concentrated 20% solution
then (10L-x) would be the amount of the 15% solution
Mixing to form a 17% 10L solution***Amount of acid in each the equalizer
.20*x + .15(10L -x) = .17*10L
solving for x
.20*x + .15*10L -.15x = .17*10L
.05x = .02*10L
x = .02*10L/.05
x = 4L, the amount of the 20% solution.
6L the amount of the 15% solution (10L-4L)
CHECKING our Answer
.20*4L + .15*6L = .8L + .9L = 1.7L = .17*10L